When calculating the cost of the trip you were using the current VAT rate of 20%,(£231-includes 20% VAT). What would the cost have been per pupil two years ago when the VAT rate was 17.5%
1
Expert's answer
2013-03-05T09:07:29-0500
When calculating the cost of the trip you were using the current VAT rate of 20%,(£231-includes 20% VAT):
£231 = £X + 0.2*£X,
X - cost of the trip without VAT
Therefore:
X = £231/1.2 = £192.5
the cost of the trip includes 17.5% VAT:
£192.5 + 0.175*£192.5 = £192*1.175= £225.6